### v1

# MobiFuse: Learning Universal Human Mobility Patterns through Cross-domain Data Fusion

Abstract
Human mobility modeling is critical for urban planning and transportation management, yet existing datasets often lack the resolution and semantic richness required for comprehensive analysis. To address this, we proposed a cross-domain data fusion framework that integrates multi-modal data of distinct nature and spatio-temporal resolution, including geographical, mobility, socio-demographic, and traffic information, to construct a privacy-preserving and semantically enriched human travel trajectory dataset. This framework is demonstrated through two case studies in Los Angeles (LA) and Egypt, where a domain adaptation algorithm ensures its transferability across diverse urban contexts. Quantitative evaluation shows that the generated synthetic dataset accurately reproduces mobility patterns observed in empirical data. Moreover, large-scale traffic simulations for the LA County based on the generated synthetic demand align well with observed traffic. On California’s I-405 corridor, the simulation yields a Mean Absolute Percentage Error of 5.85% for traffic volume and 4.36% for speed compared to Caltrans PeMS observations.
Data Fusion, Domain Adaptation, Human Mobility Pattern, Travel Trajectory
1 Introduction
1.1 Motivation
Human mobility modeling has emerged as a critical component in urban planning, management of transportation, business, and public policy development (González et al., 2008; Liao et al., 2024a; Stanford et al., 2024). Understanding how people move through cities and interact with the surrounding environment is fundamental to creating efficient, sustainable, and livable urban spaces. This importance has grown with increasing urbanization and the complexities of modern transportation systems (Hasan et al., 2013; Ma et al., 2024b).
One of the key research directions in mobility modeling lies in integrating heterogeneous data sources to address incomplete or sparse datasets. Data fusion has emerged as a critical approach for combining diverse data types and enhancing modeling accuracy and robustness. Existing methods have demonstrated the potential of integrating sources such as GPS, social media, and environmental data to provide richer insights into mobility patterns (Ounoughi & Yahia, 2023; Afyouni et al., 2022). Researchers have also explored techniques for cross-domain integration and spatial-temporal modeling to improve movement predictions and scalability (Zheng, 2015; Wei et al., 2024). Although significant progress has been made, existing approaches still encounter challenges such as data sparsity, heterogeneity, and limited adaptability across diverse geographic contexts. To address these issues, this paper proposes a context-aware data fusion framework that enhances the applicability, precision, and generalizability of mobility modeling across diverse geographic and socio-demographic settings.
Figure 1: The need for data fusion and the fused data.
1.2 Challenges
Comprehensive datasets addressing the diverse requirements of mobility research remain scarce. An ideal mobility dataset would integrate rich semantic information with high-resolution trajectory data and demographic attributes while ensuring privacy protection. However, as shown in Figure 1 (a), existing datasets often focus on only a subset of these dimensions (Kapp & Mihaljevic, 2023). The complexity of combining these diverse data sources stems from two key challenges: cross-domain data harmonization and multi-modal information integration, as shown in Figure 2.
The fundamental challenge of cross-domain data fusion lies in combining data sources that complement each other but often contain conflicting information. Our framework addresses this by fusing data listed in Figure 2, where each source has distinct limitations: trajectories lack semantic context and are usually low quality, surveys are biased and expensive, and infrastructure data requires synthesis with human mobility patterns to be meaningful.
Moreover, the integration process must also address disparities across data modalities. Spatio-temporal data requires bridging continuous trajectories with discrete locations while managing different temporal resolutions. Semantic information spans multiple representations requiring consistent interpretation, while statistical data must reconcile individual observations with aggregate patterns across appropriate spatial and temporal scales.
To address these challenges, we propose a cross-domain data fusion framework that synthesizes mobility data with integrated semantic, trajectory, demographic, and location-based information, as shown in Figure 1 (b). By leveraging the complementary strengths of these diverse data sources, our approach enables a more comprehensive understanding of mobility patterns while preserving privacy.
1.3 Contribution
Compared to prior work in data fusion and mobility modeling, we bridge multiple data gaps between: high-resolution but semantically-sparse trajectories and context-rich but limited surveys, population-level statistics and individual-level behaviors, and static infrastructure data and dynamic human movements, enabling new insights into human mobility and its implications for societal infrastructure, which hold promise for numerous applications, including traffic optimization, urban planning, and anomaly detection.(He et al., 2024; Ma et al., 2024a; Stanford et al., 2024).
Compared with the existing literature, our research makes several significant contributions:
  • We proposed a novel data fusion framework that processes and integrates cross-domain multi-modal data sources to create comprehensive synthetic datasets, while balancing usability and privacy considerations. The framework harmonizes heterogeneous data types (spatial-temporal data: GPS (Veraset, 2024) and Point of Interest (POI) (OpenStreetMap contributors, 2017), mobility data: Household Travel Survey (HTS) (Administration, 2017), demographic data: Time Use Survey (TUS) (U.S. BLS, 2024), and traffic data: road network data (OpenStreetMap contributors, 2017)) and finally validated with large-scale traffic simulation.
  • We developed a methodology that fuses the realism of GPS data with the knowledge of survey data by using semantic annotations from an LLM and reconstructing missing activities with a survey-trained model, improving the completeness and interpretability of mobility trajectories.
  • We proposed a transfer learning framework that addresses the two key data challenges: 1) enabling knowledge and pattern adaptation in data-sparse regions through semi-supervised learning, leveraging limited GPS trajectories to adapt source domain patterns without requiring complete ground truth data, and 2) incorporating demographic variations into anonymized trajectories through unsupervised distribution adaptation, aligning mobility patterns with demographic characteristics without pairwise individual labels.
Figure 2: An overview of the proposed data fusion framework. The captions within methodology indicate their corresponding sections in the paper for reference.
2 Related Work
2.1 Human Travel Trajectory Reconstruction
Trajectory reconstruction has become crucial in understanding human mobility patterns, especially when dealing with incomplete datasets. Recent advancements have addressed the challenges of data sparsity and irregularity through innovative techniques. Chen et al. (Chen et al., 2019) introduced the Context-enhanced Trajectory Reconstruction (CTR) method, using tensor factorization to reconstruct complete individual trajectories from sparse Call Detail Records. Li et al. (Li et al., 2019) proposed the Multi-criteria Data Partitioning Trajectory Reconstruction (MDP-TR) method for large-scale, low-frequency mobile phone datasets, enhancing reconstruction performance by considering spatiotemporal patterns of missing data and individual similarities. For GPS data, Zheng et al. (Zheng et al., 2010) developed a collaborative system for location and activity recommendations, demonstrating significant improvements in inferring activity types. Alexander et al. (Alexander et al., 2015) emphasized the importance of comprehensive temporal and spatial analysis in trajectory reconstruction using mobile phone data.
Despite these advancements, challenges persist, including the diverse characteristics of data sources, limited data accessibility, and restricted model adaptability across different geographic regions. These challenges underscore the need for continued research to develop more robust and widely applicable reconstruction techniques.
2.2 Large Language Models for Mobility Modeling
Large language models (LLMs) have demonstrated significant potential in mobility modeling by leveraging their ability to process contextual and semantic information from diverse data sources. Recent advancements, such as those by Liu et al. (Liu et al., 2024b) and Cai et al. (Cai et al., 2024), highlight the versatility of LLMs in tasks like trajectory reconstruction and decision-making. Studies on foundational models like LLaMA (Touvron et al., 2023), PaLM (Anil et al., 2023), and TEG-DB (Li et al., 2024) further underline their power in multi-modal data fusion and semantic interpretation. These capabilities make LLMs valuable complements to traditional methods, particularly in addressing challenges of data sparsity and complex semantic understanding in mobility applications.
2.3 Transfer Learning
Recent advancements in transfer learning have shown significant potential for enhancing human mobility analysis, particularly in data-limited scenarios. Techniques originally developed for natural language processing offer valuable insights for trajectory reconstruction and mobility modeling.
Howard and Ruder’s (Howard & Ruder, 2018) gradual unfreezing technique could be adapted to preserve general mobility patterns while adjusting to specific regional characteristics. Peters et al. (Peters et al., 2019) and Merchant et al. (Merchant et al., 2020) suggest that middle layers are often the most transferable, which could be useful when adapting mobility models from data-rich to data-scarce regions. Liu et al. (Liu et al., 2019) explored freezing bottom layers while fine-tuning top layers, a strategy that could preserve fundamental human movement patterns while adapting to unique characteristics of specific urban environments.
These studies suggest enhancing trajectory reconstruction through gradual model adaptation and selective layer fine-tuning. Future work should adapt these methods to mobility data’s unique characteristics, considering spatiotemporal dependencies and movement complexity, to improve mobility modeling in data-sparse regions.
3 Problem Formulation
To address the multi-modal and cross-domain challenges, we propose a three-stage fusion framework that advances beyond existing approaches typically limited to combining similar data sources. As shown in Figure 2, our framework integrates diverse data types to generate realistic human mobility patterns. The generative module (4.1) first employs LLMs to fuse unstructured POI data with GPS stay points, creating semantically annotated but temporally incomplete trajectories due to GPS fragmentation. These trajectories are then completed through an activity reconstruction model incorporating mobility pattern learned from travel survey.
The transfer learning stage (4.2) is designed to capture real-world mobility dynamics while enabling deployment in data-scarce regions. Through knowledge (4.2.1) and pattern (4.2.2) adaptation, the model can be adapted to different regions and cultures and learns to balance the generic knowledge learned from surveys with the realism in GPS trajectories, including region-specific patterns. Our demographic adaptation module further enhances realism by incorporating sociodemographic variations using TUS data, which provides rich activity patterns across different population groups, ensuring our synthetic trajectories reflect both demographic-specific behaviors and real-world mobility dynamics while maintaining privacy.
Finally, the third stage (4.3) grounds these behaviorally-realistic patterns in physical space by fusing the generated activity chains with spatial distributions and traffic data to assign feasible locations, enabling validation through large-scale traffic simulation. This complete integration of dynamic human behavior with static infrastructure data creates a comprehensive framework that generates mobility patterns reflecting both the complexity of human decision-making and the constraints of physical infrastructure.
Table 1: Activity category codes with descriptions
1. Home 6. Shop services 11. Social
2. Work 7. Meals out 12. Healthcare
3. School 8. Errands 13. Worship
4. Caregiving 9. Leisure 14. Other
5. Shop goods 10. Exercise 15. Pickup/Drop
Starting with LLM-based trajectory annotation, we denote $i$ for an agent. The $j$-the trajectory collected for the agent contains $N$ stay points, $Traj_{j}^{i}=\left\{P_{1}^{i,j},P_{2}^{i,j},\ldots,P_{N}^{i,j}\right\}$. Each stay point is linked with an activity, where stay point $P_{n}^{i,j}=\left[T_{n}^{i,j},(x,y)^{i,j},S_{n}^{i,j},E_{n}^{i,j}\right]$ consists of activity type $T_{n}^{i,j}$ (as shown in Table 1), GPS location $(x,y)^{i,j}$, start time $S_{n}^{i,j}$ and end time $E_{n}^{i,j}$. Then activity chains capturing the activity pattern in this region can be extracted from the annotated trajectory, as $A_{j}^{i}=\left\{A_{1}^{i,j},A_{2}^{i,j},\ldots,A_{N}^{i,j}\right\}$, where $A_{n}^{i,j}=\left[T_{n}^{i,j},S_{n}^{i,j},E_{n}^{i,j}\right]$.
With our semantically annotated trajectories represented as $A_{j}$, we propose an activity reconstruction model $M^{R1}$ tailored to a specific region $R1$. For an incomplete $A_{j}$, the model synthesis a new activity chain $A_{j}^{\prime}=M^{R1}(A_{j})$, by filling the gaps based on the common activity patterns learned from the region $R1$.
The activity reconstruction adapts cultural effects to activity patterns through transfer learning, but incorporating sociodemographic correlation patterns poses a challenge. To address this, we introduce a demographic adaptation module that modifies reconstructed activity chains based on sociodemographic characteristics. Let us define the sociodemographic space $\mathcal{S}$ as: $\mathcal{S}={\mathbf{s}_{g}|g\in\mathcal{G}}$ where $\mathcal{G}$ represents the set of demographic groups. For a demographic space with $N$ dimensions (e.g., age, gender, employment status), where each dimension $i$ has $C_{i}$ categories, the total number of demographic groups is: $|\mathcal{G}|=\prod_{i=1}^{N}C_{i}$.
The adapter functions as an additional layer after the reconstruction model $M^{R1}$. Given a batch of reconstructed activity chains probability matrix $\mathbf{Y}\in\mathbb{R}^{B\times K\times|T|}$, where $B$ is the batch size, $K=96$ represents 96 time steps with 15-minute time intervals for a day, and $|T|$ represents the number of activity types defined in Table 1, we introduce an adapter matrix $\mathbf{M}\in\mathbb{R}^{|\mathcal{G}|\times|T|}$. For each demographic group $g\in\mathcal{G}$, its adapter vector $\mathbf{a}_{g}\in\mathbb{R}^{|T|}$ modifies the reconstructed activity probabilities. The demographically-adapted activity probabilities for group $g$ across $K$ time intervals are computed as: $\mathbf{P}_{g}=\text{softmax}(\mathbf{Y}+\mathbf{a}_{g},\text{axis}=|T|)$, where $\mathbf{a}_{g}$ is broadcast across the $K$ time intervals.
Finally, we assign the activity chains to physical zonal locations while preserving spatial and temporal feasibility. Given a set of TAZs $\mathcal{Z}$ and road network $\mathcal{R}$, we assign locations to activities by optimizing:$L^{*}=\operatorname*{arg\,min}_{L\in\mathcal{Z}}\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}d(L_{i,j},L_{i,j+1})$, subject to constraints on travel time feasibility via $\mathcal{R}$ and consistency with observed spatial distributions. This enables validation against real-world traffic patterns.
4 Methodology
4.1 Generative Data Fusion
Human trajectory data collected from GPS devices include geo-location and time-stamped stay points that elucidate individual mobility patterns. However, these trajectories suffer from both semantic sparsity and temporal incompleteness. We develop the generation based data fusion module, as Figure 2 (4.1), to address these two challenges.
Adopting the methodology outlined (Liu et al., 2024a), we use LLMs to annotate stay points using the POI dataset. This process utilizes LLM to assign semantic labels to POIs based on attributes such as ”name”, ”amenity”, and ”construction type”. Each POI is then categorized into one of 15 activity codes as described in Table 1. The LLM’s prompt configuration includes a ”Task Description”, ”Activity Code Illustration”, and a description of each POI in natural language format, ensuring comprehensive understanding and accurate classification. The output from the LLM provides the top three most relevant activity codes for each POI, each with a corresponding probability.
Subsequently, a combination of rules and a Bayesian-based algorithm is used to associate stay points with the most appropriate POI and activity type, as $T=\arg\max_{T_{i}}\left\{P(T_{i}\mid\text{POI},S)\right\}$, where $T$ is the activity at the stay point, given POIs and start time $S$. This assignment takes into account both the inferred activity type and spatio-temporal information, thereby enhancing the accuracy of the trajectory’s semantic enrichment. Ultimately, this process results in human trajectories that are richly annotated with semantic information.
While this LLM-based annotation provides semantic context to the trajectories, their inherent incompleteness limits their direct use for mobility analysis. Therefore, our framework continues with an activity reconstruction process that fuses these annotated but incomplete trajectories with comprehensive mobility data from travel surveys. A base Activity Reconstruction Model $M^{0}$ is trained on survey data to learn generic activity patterns and can be adapted to other regions with small dataset, supported by the strong similarity in human activities across regions  (Schneider et al., 2013; Cao et al., 2019). The adaptation of this base model to realistic geographical and cultural contexts is detailed in Section 4.2.
Figure 3: Network architecture of the activity reconstruction model
Encoding one-day activity chains $A_{j}^{i}$ is the first step in preparing training data. This process converts a single day into a time vector with 96 time slots, each representing a 15-minute interval, where activity types occupy their corresponding time slots (see Figure 3). The training data is then processed using three masking techniques to simulate incomplete real-world data, enrich training examples, and improve the model’s robustness in representing activity patterns and interdependencies. The first technique, activity-based masking, randomly removes entire activities from the chain, simulating GPS signal loss in specific locations. The second, period masking, removes continuous periods of the day, mimicking situations where data is unavailable for extended periods, like when users are offline. The third, time slot-based masking, randomly removes individual time slots, representing sporadic or intermittent data loss.
The backbone of the model adopts Transformer-based structures (Vaswani et al., 2017) for sequence completion, enhanced with temporal features (time segments and day-of-week labels). We formulate the reconstruction as a time slot classification task with a tailored loss function that integrates three components to capture both activity accuracy and temporal dependencies:
Cross-Entropy Loss: $L_{CE}=-\sum_{c}w_{c}\cdot y_{c}\log(\hat{y}_{c})$ measures dissimilarity between predicted probabilities $\hat{y}_{c}$ and true labels $y_{c}$ with class weights $w_{c}$ to handle imbalanced classes; Transition Loss: $L_{TR}=-\frac{1}{N}\sum_{i=1}^{N}[t_{i}\log(\hat{t}_{i})+(1-t_{i})\log(1-\hat{t}_{i})]$ uses binary cross-entropy to penalize incorrect activity transitions, where $t_{i}$ and $\hat{t}_{i}$ are true and predicted transitions at time step $i$, ensuring coherent activity sequences; and Dynamic Time Warping Loss: $L_{DTW}=DTW(\hat{y},y)$ evaluates sequence similarity using DTW distance (Müller, 2007).
Finally, the loss functions $L$ combines three loss terms as $L=w_{1}\cdot L_{CE}+w_{2}\cdot L_{TR}+w_{3}\cdot L_{DTW}$.
4.2 In-context Transfer Learning
While our generation module effectively reconstructs activity chains, adapting to diverse regional contexts and demographic characteristics is crucial for realistic mobility patterns. As shown in Figure 2 (4.2), our transfer learning approach addresses this through three adaptations: 1) knowledge adaptation for region-specific knowledge when regional survey data is available, 2) mobility pattern adaptation integrating survey data with GPS data while retaining universal activity motifs, and 3) demographic-aware adaptation for demographic-specific patterns. Based on research showing that over 90% of human activities share a common structure between regions (Schneider et al., 2013; Cao et al., 2019), this underlying similarity enables efficient adaptation to new regions with limited data.
4.2.1 Supervised Transfer Learning for Knowledge Adaptation
When transfer to a region with a complete activity dataset (e.g., regional survey data), we employ progressive unfreezing as the transfer learning strategy. This approach balances knowledge preservation in the source domain with adaptation to the target domain, tuning three model components: embedding layers, Transformer layers, and an MLP, as illustrated in Figure 3. Following (Yosinski et al., 2014), input-near layers capture general features while output-near layers learn task-specific ones. By unfreezing the middle layers last (Peters et al., 2019), we mitigate catastrophic forgetting and facilitate fine-grained domain adaptation.
In particular, the unfreezing process proceeds in three phases. In the Initial Phase (first quarter of epochs), only the MLP and embedding layer are unfrozen. This allows the MLP, closest to the output, to adapt activity probability to the target domain, while the embedding layer refines semantic representations by incorporating new features. During the Intermediate Phase (second quarter of epochs), the Transformer’s encoder and decoder layers closest to the input are unfrozen, enabling adjustments to foundational sequence data patterns while preserving core source-domain knowledge. Finally, in the Final Phase, the middle layers of the Transformer’s encoder and decoder are unfrozen to refine intermediate-level representations, capturing more complex activity relationships and temporal dependencies.
4.2.2 Semi-Supervised Transfer Learning for Mobility Pattern Adaptation
A supervised learning strategy is impractical for transferring the model to regions with only incomplete GPS data. To address this data sparsity issue, we develop a semi-supervised learning strategy as a key module to fuse survey data knowledge learned by the based model with the realism of GPS data in the target region.
As outlined in Algorithm 3, the base model $M^{0}$ first synthesizes the initial dataset $A_{0}$, initiating an iterative adaptation process. At each step, synthetic datasets $A_{n}$ are generated by $M^{n}$ from raw activity chains $A_{raw}$ to train subsequent models $M^{n+1}$. The process continues until the JSD between synthetic and real-world data converges, ensuring the model’s readiness for accurate activity chain reconstruction and application in new regions.
To avoid overfitting to the new synthetic data and ensure robustness, 20% of the previous iteration’s dataset is retained in each new training cycle. This retention strategy allows the model to refine its understanding by integrating information from both new synthetic data and a portion of the previously learned data, promoting stability and preventing drastic shifts in learned patterns. Furthermore, we differentiate real and synthetic data during training by applying a mask-based loss weighting technique. This method prioritizes real data in the loss computation, weighting it more heavily in the cross-entropy loss term as: $L_{real}=\frac{1}{N_{r}}\sum_{i=1}^{N}m_{i}\cdot L_{CE_{i}}$ and $L_{synthetic}=\frac{1}{N_{s}}\sum_{i=1}^{N}(1-m_{i})\cdot L_{CE_{i}}$, where $m_{i}$ is the mask (1 for real, 0 for synthetic), $N_{r}$ and $N_{s}$ are the numbers of real and synthetic points. Then the total loss function becomes: $L=w_{1}\cdot(w_{l}\cdot L_{real}+w_{s}\cdot L_{synthetic})+w_{2}\cdot L_{TR}+w_{3}\cdot L_{DTW}$.
4.2.3 Optimization Based Transfer Learning for Demographic Adaptation
Incorporating sociodemographic patterns into mobility modeling is challenging due to scarce paired GPS-demographic data. While HTS provides such paired data, its limited regional availability constrains model generalizability. Our transfer learning approach addresses this by incorporating demographic dynamics through targeted adaptation of activity patterns. We leverage TUS as a privacy-preserving source of aggregated activity distributions across demographic groups. Harmonizing activity categories to classification space $\mathcal{T}$ yields target distributions $\mathbf{D}_{g}\in\mathbb{R}^{|\mathcal{T}|}$ for each demographic group $g\in\mathcal{G}$.
We propose a demographic adaptation mechanism through a learnable adapter matrix that modulates the activity reconstruction module’s output distribution. The adapter optimization is formulated as a optimization problem with reward-guided updates. We formulate a dual-criteria objective function that captures both global and local distribution alignment:$\mathcal{L}(\mathbf{P}_{g},\mathbf{D}_{g})=e^{-\alpha\cdot\text{RMSE}(\mathbf{P}_{g},\mathbf{D}_{g})}+\beta\cdot\text{KL}(\mathbf{P}_{g}\|\mathbf{D}_{g})$, where the exponential RMSE term ensures coarse distribution matching while the KL divergence term enforces fine-grained consistency between generated and target distributions. The adaptation weights are optimized through reward-weighted gradient descent, where $\mathcal{L}(\mathbf{P}_{g},\mathbf{D}_{g})$ serves as an adaptive scaling factor: $\mathbf{a}_{g}^{t+1}=\mathbf{a}_{g}^{t}+\eta\cdot\mathcal{L}(\mathbf{P}_{g},\mathbf{D}_{g})\cdot\nabla_{\mathbf{a}_{g}}\mathcal{L}$, with $\nabla_{\mathbf{a}_{g}}\mathcal{L}=(\mathbf{D}_{g}-\mathbf{P}_{g})$ providing the descent direction toward distribution matching.
To maintain numerical stability and interpretability, we impose bounded constraints on the adapter weights: $\mathbf{a}_{g}^{t+1}=\text{clip}(\mathbf{a}_{g}^{t+1},\epsilon,M)$, where $\eta$ denotes the learning rate, and $\epsilon,M$ define the feasible adaptation range. The optimization terminates upon reaching distribution convergence: $\|\mathbf{P}_{g}^{t}-\mathbf{D}_{g}\|_{2}\leq\delta\quad\forall g\in\mathcal{G}$, where $\delta$ controls the distribution matching precision. This formulation ensures stable convergence while maintaining interpretable demographic adaptation weights, which is elaborated at Algorithm 2.
4.3 Activity Location Assignment
To translate our generated activity chains into real-world use, we developed a location assignment module that merges spatial limits and traffic data, as in Figure 2 (4.3), which assigns each activity to specific TAZs through a layered approach: first fixing mandatory activities (home, work, school) based on commute patterns, then placing other activities while considering both distance spread and flow between fixed points. The process refines assignments until matching real spatial patterns that reflect regional traits. Through this assignment, our activity chains become real travel paths that can be tested against traffic data (detailed in Appendix A).
5 Experiments
We evaluate our framework through two case studies: Los Angeles (LA) County and Egypt. The LA study shows complete pipeline capabilities, while Egypt showcases transfer learning in data-scarce regions (as in Appendix C.2). Each component in our pipeline - semantic annotation, activity reconstruction, transfer learning (includes both activity pattern and demographic adaptation) - is essential in the data fusion process, as removing any component would prevent the generation of complete, demographically representative mobility patterns. We compare the output synthetic data against multiple sources, such as HTS, TUS and raw GPS points, as our baseline, which helped to reflect the disagreement on same measurement across different data sources.
5.1 Experimental Setup
Our dataset selection leverages the complementary strengths of different data sources: GPS trajectories provide large-scale mobility patterns but lack semantic context, which we complement with OpenStreetMap POI data. For activity pattern learning, we use HTS travel diary data as it contains complete activity chains, though its limited availability leads us to use the more accessible TUS data for demographic pattern modeling. This choice enhances our framework’s applicability across different regions. For validation, we select Caltrans PeMS traffic flow data and OpenStreetMap road network, as they provide comprehensive ground truth for assessing the realism of our generated mobility patterns.
We employ Jensen-Shannon Divergence (JSD) as our primary metric to evaluate distributional similarities across multiple dimensions: activity chain length, duration, type, and temporal patterns (start/end times). Lower JSD values indicate better alignment. For spatial validation, we use cosine similarity for Origin-Destination matrices and Mean Absolute Percentage Error (MAPE) for traffic metrics.
5.2 Model Performance Analysis
5.2.1 Base Model Evaluation on Survey Data
The performance of the base model $M^{N}$ in reconstructing 70% masked NHTS (National HTS) data and its transferability to the SCAG datasets is quantified in Table 2.
Table 2: JSD value for similarity measurement across datasets and model performance
JSD Values Length Duration Type Start Time End Time
NHTS vs SCAG 0.017 0.009 0.059 0.018 0.013
$M^{N}$ on NHTS 0.017 0.002 0.001 0.001 0.005
$M^{N}$ on SCAG 0.021 0.003 0.017 0.004 0.007
$M^{S}$ on SCAG 0.011 0.004 0.004 0.001 0.001
Figure 4: Performance evaluation on NHTS and SCAG with 70% masked activity chain
The model achieves strong performance on NHTS data, with JSD values ranging from 0.001 to 0.017 across all metrics. However, directly applying the model to SCAG data results in performance degradation indicated by increasing JSD value, revealing significant regional differences in mobility patterns. The JSD values between NHTS and SCAG, which range from 0.009 to 0.059 across various activity attributes, quantify these differences and are visualized in Figure 4. Supervised transfer learning is applied to $M^{N}$ to obtain $M^{S}$, yielding low JSD values when reconstructing the SCAG dataset. This result shows the algorithm effectively adapts to different geographical contexts and data formats, despite initial disparities between SCAG and NHTS datasets.
5.2.2 Transfer Learning Evaluation on GPS and Survey Data
JSD values across iterations for the evaluation metrics are shown in Table 3, showing how the iterative transfer learning process improves alignment with the target region.
Table 3: Model evolution during iterative transfer learning
JSD Values Activity Type Start Time End Time
NHTS vs LA GPS 0.0508 0.0289 0.0367
Iteration_1 0.0116 0.0014 0.0037
Iteration_2 0.0033 0.0013 0.0038
Iteration_3 0.0022 0.0008 0.0027
Iteration_4 0.0025 0.0007 0.0019
Iteration_5 0.0046 0.0024 0.0052
Iteration_6 0.0077 0.0048 0.0081
In this table, only three metrics are included because the lack of ground truth in incomplete data renders activity duration and chain length unreliable for comparison, whereas the selected metrics are more accurate at the system level and can better reflect regional mobility patterns. Activity type JSD drops from 0.0508 at baseline to 0.0022 by Iteration 3. Similarly, start time and end time JSDs reach their lowest values, 0.0007 and 0.0019 respectively, at Iteration 4. Model performance deteriorates beyond Iterations 3 and 4, which appear to strike the best balance for reducing dissimilarities and adapting the model. This decline aligns with findings that AI models collapse when trained on recursively generated data (Shumailov et al., 2024), as over-reliance on synthetic data erodes nuanced or rare activities in the original data.
We evaluate our demographic transfer learning module by first comparing generated activity patterns across demographic groups with both TUS and HTS data (Table 4). Note that these JSDs measure alignment at the demographic group level, where matching patterns becomes inherently more challenging than population-level alignment due to the need to capture group-specific behavioral variations. The results demonstrate effective demographic adaptation, with output patterns showing strong alignment with TUS data (JSD: 0.006-0.018) across all demographic groups. This close alignment with TUS validates our adapter’s ability to learn demographic-specific patterns. The larger differences between output and HTS patterns (JSD: 0.048-0.094) are expected and reasonable, as they reflect the inherent differences between TUS and HTS data (TUS vs HTS JSD: 0.011-0.063). This suggests our framework successfully adapts to TUS patterns while maintaining reasonable consistency with HTS characteristics.
Table 4: JSD value for similarity measurement of activity duration distribution across demographic group
Group HTS vs Output TUS vs Output TUS vs HTS
Young M .083/.086 .008/.009 .014/.014
Young F .094/.085 .015/.013 .016/.011
Mid-aged M .048/.067 .018/.006 .016/.054
Mid-aged F .055/.066 .018/.007 .022/.059
Elder M .070/.066 .015/.009 .020/.055
Elder F .077/.071 .013/.009 .029/.063
Note: Values shown as (Worker/Non-worker); M: Male F: Female
Beyond activity durations, we evaluate temporal patterns across multiple scales: demographic groups (Figure 5), aggregate patterns (Figure 10), and activity-specific distributions (Figure 11). For demographic groups, the results show consistent patterns for mid-aged and elderly groups, while young groups show discrepancies due to limited young worker representation in HTS (2% of population). At the aggregate level, the model achieves strong alignment with both GPS data (JSD: 0.004) and NHTS patterns (JSD: 0.014), capturing key temporal features including peak periods. Major activities maintain their characteristic patterns through the transfer learning process, while our framework also captures early morning activities (before 6 AM) missing from HTS data due to survey limitations like underreporting and rounding effects (Bricka et al., 2012). This validates our approach’s ability to balance GPS and survey data while preserving demographic characteristics.
Figure 5: Temporal performance evaluation per demographic group based. Note: HTS lacks a Young Worker group; comparison for young individuals ignores employment status.
5.3 Traffic Simulation and Multi-Scale Validation of Synthetic Trajectories
To validate that our generated activity chains reflect realistic mobility patterns, we generated 1 million synthetic activity chains based on above methodology, and using location assignment module to assign perform a comprehensive real-world validation through large-scale traffic simulation, using MATSim (Horni et al., 2016) across LA County’s road network, enabling comparison with Caltrans traffic data (CalTrans, 2020). Our multi-scale validation shows strong performance across all spatial levels: at the LA regional level, the model achieves 0.997 cosine similarity for origin-destination matrices across eight sub-regions, demonstrating accurate reproduction of macro-level travel patterns and spatial interaction between regions; at the system level, MATSim simulations achieve MAPE of 4.97% for Vehicle Miles Traveled (VMT) and 1.16% for traffic speed, indicating that our synthetic populations generate realistic system-level traffic behavior across the entire network; at the corridor level, analysis of the I-405 corridor shows MAPE of 5.85-9.32% for traffic volume and 4.36-4.45% for speed across different directions, capturing local patterns including peak-hour congestion, validating the model’s ability to capture fine-grained local traffic patterns including peak-hour congestion. These results validate our framework’s ability to generate human travel diary with realistic mobility patterns, as evidenced by strong agreement with real-world traffic data (detailed in Appendix C.1.3).
6 Conclusion
This paper presents a framework that enables realistic mobility pattern generation through data fusion. Our approach addresses fundamental data challenges: semantically enriching and completing trajectories through LLM annotation and reconstruction, adapting to new regions through transfer learning, and incorporating demographic patterns without paired labels. We validate the generation based on real-world test case throughout large-scale simulation and showcase the strong performance. By fusing diverse data sources, our framework demonstrates the capability of overcoming individual data limitations. Future work could extend this framework to capture more complex social dynamics, including coordinated activities between individuals, large gathering events, and social network effects on mobility patterns, further enhancing the realism of generated trajectories.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning (ML), potentially benefiting transportation efficiency, urban development, and sustainability efforts. Our framework demonstrates how ML can harmonize diverse data sources to generate valuable insights while maintaining privacy. All data sources used are either publicly available (OpenStreetMap, census data, travel surveys) or commercially available in pre-anonymized form (GPS trajectories), with no collection of personal identifiers or sensitive information. The framework’s ability to work with anonymized data and generate synthetic patterns further ensures privacy preservation while enabling broad applicability for urban planning and policy making.
Appendix A Detailed Methodology of Activity Location Assignment
The activity location assignment process transforms abstract activity chains into spatially-grounded trajectories by assigning each activity to specific TAZs. This process must balance multiple objectives: maintaining realistic commuting patterns, preserving feasible trip chains, and matching observed spatial distributions of activities across the region.
A.0.1 Mandatory Activity Assignment
The first stage focuses on assigning locations for mandatory activities (work and school), using home locations as fixed anchor points. This stage is crucial as these locations strongly influence the spatial patterns of other activities throughout the day. The process includes:
1. Distance Sampling: For each individual, we sample commute distances from demographic-specific distributions ($D_{md}$) derived from regional data. These distributions vary by sub-region to capture local commuting patterns, such as longer commutes in suburban areas versus shorter ones in urban cores.
2. Land Use Compatibility: Candidate TAZs are filtered based on land use compatibility. For work activities, we consider employment density and workplace location patterns; for school activities, we account for educational facility distribution.
3. Spatial Matching: The final TAZ selection optimizes for both the sampled distance and land-use suitability while ensuring the spatial distribution of mandatory activities matches regional patterns. This is formulated as: $Z_{md}=\operatorname*{arg\,min}_{z\in\mathcal{Z}}\{d(h,z)-\hat{d}\}$ where $h$ is the home TAZ, $z$ is a candidate TAZ, and $\hat{d}$ is the sampled target distance.
A.0.2 Non-mandatory Activity Assignment
The second stage assigns locations for non-mandatory activities (shopping, leisure, etc.) between mandatory activity anchor points. This stage employs a more complex optimization approach:
1. Distance and Angular Constraints: Two key parameters guide the assignment: - Distance to next activity ($D_{nmd}$): sampled from activity-specific distributions - Angular deviation ($\theta$): measures the difference between the direct path to the next activity and the path to the next anchor point
2. Trip Chaining Optimization: Location selection considers both previous and next activities through a weighted objective function: $Z_{nmd}=\operatorname*{arg\,min}_{z\in\mathcal{Z}}\{\alpha\cdot|d(z_{prev},z)-\hat{d}|+\beta\cdot|\theta(z_{prev},z,z_{next})-\hat{\theta}|\}$ where $\alpha$ and $\beta$ balance the importance of distance and directional constraints.
3. Temporal Feasibility: Assignments must ensure travel times between activities are realistic given the transportation network. This adds a constraint: $t(z_{i},z_{i+1})\leq T_{max}$ where $t(\cdot)$ is the travel time function and $T_{max}$ is the available time window.
A.0.3 Distribution Refinement
The final stage ensures that the aggregate spatial distribution of activities matches observed patterns through iterative refinement:
1. Distribution Comparison: Activity frequencies across sub-regions are compared with ground truth data using cosine similarity metrics.
2. Parameter Adjustment: Reference distributions ($D_{md}$, $D_{nmd}$, $\theta$) are iteratively adjusted using: $D^{t+1}=D^{t}+\eta\cdot(F_{target}-F_{current})$ where $F$ represents activity frequency distributions and $\eta$ is a learning rate.
3. Convergence Criteria: The refinement process continues until: $|F_{target}-F_{current}|<\epsilon$ where $\epsilon$ is a predefined threshold.
The complete algorithmic implementation is presented in Algorithm 1, which formalizes this three-stage process into a computational framework. The algorithm employs efficient data structures and optimization techniques to handle large-scale assignment tasks, making it suitable for regional-level applications.
Algorithm 1 Location Assignment for Activity Chains
1:  Input: Activity chains $A$, Home TAZ assignments $H$, Reference patterns $R$
2:  Output: Spatially-assigned activity chains $A^{\prime}$
3:  for each activity chain $a_{i}\in A$ do
4:     $h_{i}\leftarrow H[i]$Get home TAZ}
5:     // Assign mandatory activities
[…]
7:        $d\sim P(D|\text{type}(k))$Sample from distance distribution}
8:        $k.\text{taz}\leftarrow\operatorname*{arg\,min}_{z}|d(h_{i},z)-d|$
9:     end for
10:     // Assign non-mandatory activities
11:     for activity $k\in a_{i}$ where $\text{type}(k)$ is non-mandatory do
12:        $p\leftarrow\text{prev\_anchor}(k)$
13:        $n\leftarrow\text{next\_anchor}(k)$
14:        $d\sim P(D|\text{type}(k),\text{region}(h_{i}))$
15:        $\theta\sim P(\Theta|\text{type}(k))$
16:        $k.\text{taz}\leftarrow\operatorname*{arg\,min}_{z}\{\alpha|d(p,z)-d|+\beta|\theta(p,z,n)-\theta|\}$
17:     end for
18:  end for
19:  return $A^{\prime}$
This location assignment methodology ensures that the generated activity patterns are not only temporally and demographically realistic but also spatially feasible within the context of real urban environments. The multi-stage approach allows for fine-grained control over spatial distributions while maintaining the integrity of individual activity chains.
Appendix B Dataset Details
B.1 Household Travel Surveys
Household Travel Surveys (HTS) have long been a fundamental data source in transportation planning and modeling. These surveys provide comprehensive insights into daily travel patterns by collecting detailed information about individual and household travel behaviors, including complete door-to-door trip chains, activity purposes, timing, and mode choices. This detailed trip chain information makes HTS particularly valuable for understanding human mobility decisions’ sequential nature and underlying patterns.
The training of the mobility pattern reconstruction model requires the use of complete activity chain, including every activity of an agent in a 24-hour day. Two datasets with completed activity chains were utilized in this study to construct the base model.
In order to learn common mobility patterns at the national level, we used the 2017 National Household Travel Survey (NHTS), administered by the Federal Highway Administration (FHWA), to capture nationwide mobility patterns (Administration, 2017). The NHTS includes socio-economic demographics and travel diaries from 129,600 households. We aggregated the original 19 activity types recorded in the NHTS into 15 categories based on the activity locations, which are presented in Table 1. In this study, only the travel diary of each agent in NHTS is used for training.
To learn a specific regional activity pattern, we adopted the activity chain data generated by the Activity-Based Model (ABM) of Southern California Association of Governments (SCAG). The SCAG ABM is a comprehensive modeling system that integrates a series of activity-related choice models, ranging from long-term to short-term decisions, to simulate the activity chains of residents in Southern California. This model generates a synthetic dataset that includes detailed information about the types and sequences of activities, locations, and start/end times. These features make the dataset a valuable tool for analyzing the spatiotemporal dynamics of human activities. Further details of the SCAG ABM and its methodology can be found in  (He et al., 2022; Jiang et al., 2022).
B.2 Time Use Surveys
Time Use Surveys (TUS) are systematic studies that collect detailed information about how individuals spend their time throughout the day. These surveys typically require participants to maintain detailed diaries, recording their activities in fixed time intervals (often 15-30 minutes), including information about what they do, when they do it, where they are, and who they are with. Unlike transportation-focused surveys, TUS captures the full spectrum of human activities, from personal care and work to leisure and social interactions.
A key advantage of TUS is its widespread availability across different countries and regions. National statistical offices regularly conduct these surveys as part of their standard data collection efforts, making them more accessible than specialized transportation surveys like HTS. This broad availability is particularly valuable for regions with no HTS. The generalizability of our framework partially comes from the usage of TUS, which can replace the generation’s reliance on HTS with a more accessible data source.
Time Use Surveys (TUS) contain detailed individual-level records of daily activities, where each respondent reports their time allocation alongside their demographic attributes. We leverage this data in two ways: first, the individual records with demographic labels are used to train a demographic classification model, enabling automatic categorization of semantic enriched GPS trajectories into demographic groups. Second, by aggregating these individual records by demographic attributes, we obtain target activity distributions for each demographic group. These aggregated distributions then serve as optimization targets for our demographic adapters optimized by Gradient-Based Distribution (Algorithm 2), allowing the model to adjust activity patterns according to the classified demographic characteristics.
Algorithm 2 Gradient-Based Distribution Optimization for Demographic Adapters
  Input: Reconstruction Model $f$, Time Use Survey Data $\mathcal{T}$, Learning rate $\eta$, Convergence threshold $\delta$
  Output: Optimized adapter collection $\mathcal{A}$
  Process $\mathcal{T}$ to get target distributions $\{\mathbf{D}_{g}\}$
  Initialize $\mathcal{A}=\{\}$
  for $g\in\mathcal{G}$ do
     Initialize adapter weights $\mathbf{a}_{g}$
     repeat
        $\mathbf{Y}=f(\text{input})$
        $\mathbf{P}_{g}^{t}=\text{softmax}(\mathbf{Y}+\mathbf{a}_{g}^{t})$
        $\mathcal{L}=e^{-\alpha\cdot\text{RMSE}(\mathbf{P}_{g}^{t},\mathbf{D}_{g})}+\beta\cdot\text{KL}(\mathbf{P}_{g}^{t}|\mathbf{D}_{g})$
        $\nabla_{\mathbf{a}_{g}}\mathcal{L}=\mathbf{D}_{g}-\mathbf{P}_{g}^{t}$
        $\mathbf{a}_{g}^{t+1}=\mathbf{a}_{g}^{t}+\eta\cdot\mathcal{L}\cdot\nabla_{\mathbf{a}_{g}}\mathcal{L}$
        $\mathbf{a}_{g}^{t+1}=\text{clip}(\mathbf{a}_{g}^{t+1},\epsilon,M)$
        if $\|\mathbf{P}_{g}^{t}-\mathbf{D}_{g}\|_{2}\leq\delta$ then
           break
        end if
     until $iteration=max\_iterations$
     $\mathcal{A}=\mathcal{A}\cup\{\mathbf{a}_{g}\}$
  end for
  return $\mathcal{A}$
B.3 GPS and POI Data Integration
GPS trajectory data provides rich spatiotemporal information about human movement patterns. In our framework, we utilize the Veraset dataset, which contains anonymized cellular points and their trajectories. While this data offers invaluable insights into real-world movement patterns, it comes with several inherent limitations. First, GPS data often contains measurement errors and noise, leading to inconsistent or inaccurate location recordings. More significantly, due to privacy concerns, the data is anonymized, stripping away crucial information about trip purposes and individual characteristics. This anonymization, while essential for protecting personal privacy, creates a significant gap in understanding the underlying motivations for observed movements.
These limitations of raw GPS data can be largely addressed through strategic integration with POI data. POI data provides the semantic context necessary to transform anonymous coordinate points into meaningful activity locations. By matching GPS stay points with nearby POIs, we can infer likely activity types and purposes, adding crucial semantic meaning to the raw movement data. Our framework employs Large Language Models (LLMs) for this matching process, enabling more nuanced and context-aware association between locations and activities.
However, the integration of GPS and POI data presents its own challenges. POI databases, particularly from open-source platforms, often suffer from completeness and accuracy issues. Some locations may be missing, incorrectly categorized, or outdated. Additionally, in dense urban areas, multiple POIs might exist within the error radius of a GPS stay point, making it difficult to determine the actual visited location. Our framework addresses these challenges through a probabilistic matching approach that considers both spatial proximity and temporal patterns, supplemented by the semantic understanding capabilities of LLMs.
The complementary nature of these data sources enables our framework to generate more realistic and contextually aware mobility patterns while maintaining individual privacy and data security. This integration forms a crucial foundation for subsequent stages of our mobility modeling pipeline, particularly in activity reconstruction and location assignment.
Appendix C Experiment Details
C.1 Los Angeles County Case Study
C.1.1 Data Preparation
Figure 6: Example POI labeling (Liu et al., 2024a)
Figure 7: Example annotated trajectory (Liu et al., 2024a)
To begin our first case studies of LA County, we gathered half-years of GPS points from Veraset, a data vendor. We preprocess the raw GPS trajectories to daily stay points data aggregated by agents, which are consecutive minimum 10-minutes dwell points throughout the GPS trajectory. We then applied filters to acquire around 1 million of relatively high-quality daily stay points data, with a minimum coverage of 55% time of day and more than 1 activity per day per stay points trajectory. Then we fuse the daily stay points and POI data to annotated activity chains using the LLM-based semantic enrichment approach developed in (Liu et al., 2024a). This method combines POI classification using Large Language Models with a Bayesian-based activity inference algorithm to identify the most probable activity type for each stay point. Examples of POI labeling and annotated trajectories are shown in Figure 6 and Figure 7.
Then we construct a representative synthetic population for Los Angeles County using a rule-based classification approach. Through analysis of our annotated trajectories, approximately 70% contain work-related activities, aligning with LA County’s employment rate (US Bureau of Labor Statistics, 2023). We leverage distinctive activity patterns as classification rules - trajectories containing work activities are classified as belonging to workers, while those with school activities are strongly indicative of young demographics. This rule-based approach, rather than learning-based methods, was deliberately chosen to minimize error propagation in upstream processing stages that could compound into larger biases in downstream tasks. For trajectories where simple rules cannot provide confident classification, we maintain them as unclassified rather than risk misattribution.
C.1.2 Base Model and Its Transferability
The above process prepares semantically enriched but incomplete activity chains for our core reconstruction task. For base model training, we utilize the Southern California Association of Governments (SCAG) Activity-Based Model (ABM) dataset, which provides comprehensive synthetic activity chains for the Los Angeles region. The SCAG ABM integrates various activity-related choice models spanning both long-term and short-term decisions, generating detailed information about activity types, sequences, and temporal patterns (Liao et al., 2024b). We selected this dataset over alternatives like NHTS due to its substantial size - 700,000 samples for training, 200,000 for validation, and 100,000 for testing - which is crucial for effectively training transformer-based architectures.
The training procedure follows a carefully structured three-phase protocol: 1. Initial Warm-up Phase: The model trains on unmasked data for 5 epochs, establishing baseline understanding of complete activity chains 2. Intermediate Phase: Training continues with 40% masked data for 40 epochs, introducing the model to partial data scenarios 3. Final Phase: The remaining epochs use 70% masked activity chains, preparing the model for heavily incomplete real-world data
To ensure robust training, we implement several regularization techniques including dropout, L2 regularization, and early stopping. The model is trained for 120 epochs with a batch size of 512, optimizing a combined loss function that incorporates cross-entropy, transition, and dynamic time warping losses.
The model’s performance in reconstructing activity chains is first evaluated on two survey datasets with ground truth data. Using the test set, 70% of the activity chains are masked and reconstructed. The evaluation measures the similarity between the reconstructed and original activities based on daily activity counts (chain length), activity durations, types, and start/end times.
Algorithm 3 Semi-Supervised Transfer Learning
  Input: Raw activity chain $A_{raw}$, initial model $M^{0}$
  Output: Synthetic activity chain $A_{syn}$
  Generate $A_{0}=M^{0}(A_{raw})$ and initialize $n=1$
  repeat
    Train $M^{n}$ using $A_{n-1}$ with masking
    Generate $A_{n}=M^{n}(A_{raw})$
    Compute JSD similarity $J=\text{JSD}(A_{raw},A_{n})$
    if $J$ has converged then
       $A_{syn}=A_{n}$
       break
    end if
    $n=n+1$
  until convergence or maximum iterations reached
The good performance of the base model $M^{N}$ (trained on NHTS data) is quantified by its values reconstructing masked NHTS data, achieving JSD values between 0.001 and 0.017 in Table 2. However, when applied to the SCAG dataset without transfer learning, the model’s performance declines, with significant JSD values increasing. This indicates that the model trained on national-wide data doesn’t adapt well to regional (Los Angeles) patterns, due to the mobility pattern differences between these two datasets. The JSD values between NHTS and SCAG, ranging from 0.009 to 0.059 across various activity attributes, quantify these differences. For instance, as shown in Figure 4, the activity start time distribution shows a notable peak for SCAG (in red) around 8:00 am, which is less pronounced in the NHTS data (in blue). Moreover, the activity type graph reveals that SCAG data lacks representation in categories like childcare, health care visit, and religious activities, which are present in the NHTS dataset.
To address these discrepancies, we implement transfer learning on model $M^{N}$ (algorithm 3. This approach yields marked improvements, with the SCAG dataset fine-tuned model $M^{S}$ achieving JSD values as low as 0.001 for time and 0.001 for activity type. Figure 4 visually confirms this improvement, showing close alignment between red solid lines and red dash lines across all distributions. These results demonstrate our algorithm’s effectiveness in adapting to different geographical contexts and capturing diverse mobility patterns, despite the initial significant disparities between the SCAG and NHTS datasets.
Figure 8: The start time of common activities reconstructed by base and its transferred model using NHTS and SCAG
Beyond its commendable performance in system-level JSD evaluations, the proposed model exhibits proficiency in capturing patterns of common activities, with the transfer learning successfully adapting the model from regional to national contexts. As illustrated in Figure 8, the model demonstrates remarkable accuracy in reconstructing the temporal patterns of activity start times across both datasets.
Essential activities such as home, work, and school show closely aligned patterns between the two datasets, evidenced by low JSD values (0.023, 0.021, and 0.024 respectively), highlighting the universality of these routines. Conversely, regional variations are apparent in other activities. Shopping patterns, for example, differ significantly, with SCAG data showing a midday peak (JSD 0.054) versus NHTS’s more even distribution, suggesting distinct regional shopping behaviors. Exercise activities also vary, with NHTS indicating a preference for early morning sessions (JSD 0.036). These differences likely mirror varying lifestyle choices between urban Los Angeles and the broader national context. The model’s consistent low $M^{S}$ and $M^{N}$ values, typically below 0.01, confirm its efficacy in capturing these nuances, underscoring its utility for urban planning and policy-making by identifying both universal and regional behavioral patterns.
The JSD values for the evaluation metrics across iterations are presented in Table 3. Only three metrics are evaluated, because activity duration and activity chain length lack ground truth in incomplete data, making them unreliable for comparison. In contrast, the distributions of activity type, start time, and end time are more accurate at the system level and effectively reflect regional mobility patterns. The Table 3 shows the iterative transfer learning process improving alignment with the target region. Activity type JSD decreases significantly from 0.0508 (baseline) to a minimum of 0.0022 at Iteration 3, demonstrating effective adaptation. Start time and end time JSDs reach their lowest values (0.0007 and 0.0019, respectively) at Iteration 4. However, slight increases in later iterations suggest diminishing returns or potential overfitting. Overall, the results highlight that Iterations 3 and 4 are optimal for reducing dissimilarities across these metrics, reflecting successful model adaptation. Figure 9 demonstrates how the progressive adaptation of the NHTS-based model to the LA context over multiple iterations via the proposed Semi-Supervised Transfer Learning using LA GPS data.
Figure 9: Iterative adaptation from NHTS to LA via semi-supervised transfer learning with LA GPS data
The second step of transfer learning is performing the demographic adaptation by fusing TUS data. We trained 12 demographic adapters in accordance with sex (male/female), age (young, mid-aged, elderly), and employment status (worker, non-worker). To establish target distributions for each demographic group, we aggregate TUS data to create activity pattern benchmarks. Analysis of the aggregated Time Use Survey distributions (Table 5) reveals distinct activity patterns across demographic groups with aspects of sex, age, and employment status. We proceed with adapter training for each demographic group. The training process demonstrates varying levels of convergence and efficiency across different demographic segments. Overall, the adaptation process shows strong performance, with reward values (measuring similarity to target distributions) consistently above 0.85 across all demographic groups, and many groups achieving rewards above 0.95.
The training patterns reveal interesting demographic-based variations in adaptation difficulty. Worker groups generally achieve higher reward values (0.95-0.98) and converge more quickly, typically within 400-500 epochs for Mid-aged and elderly workers. This faster convergence likely reflects more structured and consistent daily patterns among employed individuals. Young workers require more training epochs ($>1000$) but ultimately achieve similarly high reward values, suggesting more complex activity patterns that require additional learning time. Non-worker groups present greater challenges in pattern adaptation, reflected in slightly lower reward values (0.82-0.93) and longer training times. This increased difficulty in modeling non-worker patterns aligns with the greater variability in their daily activities observed in the target distributions.
Table 5: L.A. County major activity duration (minutes/day) comparison with JSD
Demographic Group (JSD) Activity Duration (minutes)
Home Work School Buy goods Buy meals Recreational Exercise Visit friends
(1) (2) (3) (5) (7) (9) (10) (11)
Young Male W. (0.0085) 843.71/837.50 209.58/224.50 117.76/134.74 19.61/15.01 52.49/38.51 100.71/53.32 23.82/18.55 61.53/53.38
Young Female W. (0.0159) 820.18/785.27 176.23/204.68 131.50/134.76 39.82/20.84 57.31/34.26 70.19/34.58 23.50/18.79 75.84/88.82
Mid-Aged Male W. (0.0189) 798.36/800.37 380.79/382.87 0.00/0.00 37.89/19.10 61.70/76.96 76.16/62.16 23.92/4.67 42.35/0.98
Mid-Aged Female W. (0.0185) 839.04/825.56 332.97/330.77 0.00/0.00 42.36/22.70 57.00/74.31 63.22/58.69 16.75/4.43 46.14/2.07
Elderly Male W. (0.0157) 875.00/886.00 295.59/284.50 0.00/0.00 34.75/29.95 69.24/77.71 83.23/61.21 20.51/4.41 35.58/0.80
Elderly Female W. (0.0136) 913.23/902.44 245.65/260.23 0.00/0.00 41.41/35.62 62.34/67.56 84.74/60.80 15.03/3.99 36.99/0.83
Young Male N. (0.0090) 903.82/960.43 0.00/0.00 170.74/203.06 21.34/14.47 54.04/68.70 97.69/84.56 57.86/41.07 52.06/44.74
Young Female N. (0.0103) 899.88/987.70 0.00/0.00 170.54/149.18 36.01/21.77 59.85/75.15 79.72/76.87 31.51/39.91 49.20/40.00
Mid-Aged Male N. (0.0064) 1013.83/1054.20 0.00/0.00 18.49/3.80 40.56/42.97 57.17/69.00 127.28/118.24 26.97/37.15 45.55/32.73
Mid-Aged Female N. (0.0078) 993.96/1083.50 0.00/0.00 11.88/3.86 44.16/22.48 58.29/64.68 90.56/67.73 16.93/11.26 54.37/57.17
Elderly Male N. (0.0092) 1022.03/1101.41 0.00/0.00 1.56/3.34 42.11/43.75 70.30/80.47 144.36/150.71 25.04/5.76 41.62/22.40
Elderly Female N. (0.0099) 1037.10/1115.56 0.00/0.00 1.63/3.32 37.40/41.37 63.45/88.91 120.76/128.88 15.08/4.74 46.89/31.46
Note: Values shown as TUS/Output. W: Worker, N: Non-worker. JSD values indicate distribution similarity (lower is better).
The entire adapter training process, implemented on an NVIDIA RTX 4090 GPU, requires approximately 10 hours of computation time, averaging 40 minutes per demographic group. This relatively efficient training time, despite the complexity of demographic-specific pattern learning, demonstrates the practical feasibility of our approach for real-world applications.
After training the demographic-aware adapters, we integrate them with the transferred reconstruction model to create a comprehensive system capable of generating both demographically appropriate and temporally coherent activity patterns. This integrated model preserves the base model’s capabilities in temporal pattern recognition and activity dependencies while incorporating demographic sensitivity through the trained adapters.
Using the trained demographic-aware model, we conduct large-scale inference to generate synthetic activity patterns. We handle different trajectory data using different masking methods. For trajectories with relatively low loss or high completeness, we can often infer their socio-demographic information. In such cases, the problem becomes one of completion: we simply match the corresponding socio-demographic info to the trajectory and use the model to fill in the missing data. However, for data with low completeness, it is not possible to infer any socio-demographic information. We treat this as a generation task, masking as much data as possible—up to 80% of the time in a day—and assigning it a socio-demographic profile. The model then generates a trajectory consistent with this socio-demographic information.
Figure 10: Comparison of aggregated activity start time distributions across data sources (NHTS Survey, Veraset GPS Data, and Model-Generated Data)
The model generates activity chains through multiple passes - 5 passes for weekdays and 2 passes for weekends - with each pass using different random masks to ensure diversity in the generated patterns. This process allows the model to explore various possible activity combinations while maintaining demographic consistency. The generation process incorporates both demographic factors and variations in the day of the week to ensure realistic activity patterns between different segments of the population and throughout the week.
The generation demonstrates strong alignment with Time Use Survey across all demographic groups. Quantitative evaluation using JSD shows excellent performance (detailed results in Table 5), with non-worker groups achieving particularly strong alignment (JSD 0.0064-0.0103) and worker groups showing comparable performance (JSD 0.0085-0.0189). Mid-aged male non-workers achieve the best performance (JSD = 0.0064), while maintaining realistic activity patterns such as work duration differences across age groups and gender-specific activity preferences.
Figure 11: Activity-specific temporal pattern comparison: start time distributions for major activities across different data sources
To validate our transfer learning fusing module successfully learns the activity and demographic nuances from multi-source data, we analyze start time distributions across major activity types, comparing our generated patterns with both NHTS and raw GPS input data. The analysis is conducted at an aggregate level rather than per demographic group to evaluate the model’s overall capability in preserving temporal dynamics and to provide sufficient sample sizes for meaningful comparison, particularly for less frequent activities. We visualize and quantify these comparisons through probability distributions and JSD metrics to assess both pattern similarity and statistical significance.
As shown in Figure 10, the aggregated activity start time distribution demonstrates strong alignment across all other data sources. The generated patterns show particularly close agreement with the raw GPS data (JSD: 0.004), confirming the model’s ability to preserve real-world temporal patterns while incorporating demographic variations. This corresponds to findings from (Stopher & Greaves, 2007), which highlight the temporal precision of GPS data compared to traditional survey methods. The comparison with NHTS data also shows strong correspondence (JSD: 0.014), with matching morning (8:00–10:00) and afternoon (16:00–18:00) peak periods, though with slight probability differences of approximately 0.005. These discrepancies reflect the systematic differences between survey and GPS data collection methods, such as recall bias and rounding effects, where respondents often report activities at common clock hours (Bricka et al., 2012), while GPS data captures actual movement patterns with higher temporal precision. At the individual activity level, as shown in Figure 11, the model maintains high fidelity across both major mandatory and non-mandatory activities, with average JSD values consistently below 0.06. We noticed that the generation distribution shapes strike a balance between the disagreement of GPS and Survey data and create a reasonable temporal activity patterns credit to the robust learning capabilities of our fusion and transfer learning pipeline.
C.1.3 Spatial Level Validation
We utilize the generated output from prior pipeline to assign locations in the context of LA County using Algorithm 1, with a spatial division of eight sub-regions (Figure 12(a)). The model is first calibrated using 100,000 agents from the SCAG Activity-Based Model (ABM) dataset, then scaled to generate mobility patterns for 1 million agents.
Figure 12: LA County Map and Freeway Network.
Regional-level Analysis: The spatial distribution of activities across sub-regions demonstrates the effectiveness of our location assignment method (Figure 12(a)). The high cosine similarity (0.997) between the origin-destination matrices of our model and SCAG ABM validates the preservation of regional flow patterns.
Figure 13: Validation for Location Assignment and traffic loading at network level.
System-wide Traffic Performance: Traffic simulation on the LA transportation network (Figure 12(b)) reveals strong system-level performance. The hourly VMT and traffic speed patterns closely match the benchmark SCAG ABM results, with MAPEs of 4.97 and 1.16 respectively (Figure 13(b)). These metrics indicate accurate reproduction of aggregate travel patterns.
Figure 14: Validation for Location Assignment and traffic loading at corridor level.
Corridor-level Analysis: We conducted detailed analysis on a major segment of Interstate 405, comparing our results with both SCAG ABM predictions and PeMS real-world observations. The analysis reveals:
  • Directional variations in congestion patterns are well-captured
  • Northbound traffic: MAPE of 5.85 for volume, 4.45 for speed
  • Southbound traffic: MAPE of 9.32 for volume, 4.36 for speed
The corridor-level results (Figure 14) demonstrate the model’s capability to reproduce fine-grained traffic dynamics, including peak-hour congestion and directional flow patterns.
C.2 Egypt Case Study
C.2.1 Regional Adaptation in Data-Scarce Contexts
To evaluate the regional adaptability of our proposed framework, we conducted a case study on Egypt, a data-scarce region. While the Los Angeles case demonstrated the framework’s ability to reflect demographic and temporal characteristics of human mobility, the Egypt case emphasizes its capability to adapt to new cultural and regional contexts. We utilized a dataset of 479,526 samples, rigorously curated to ensure quality by retaining only agents with at least 7 days of recorded activity and sufficient daily temporal coverage. This dataset was partitioned into training (80%) and validation (20%) sets.
Demographic information was assigned to the dataset based on Egypt’s demographic distribution, incorporating two dimensions: age and sex. Each agent was categorized into one of six demographic groups: young, mid-aged, or elderly males and females. The GPS trajectory annotation follows the same methodology. Notably, the LLM demonstrated robust cross-linguistic capabilities, accurately matching Arabic-language POIs with activity types. This enabled the integration of diverse POI data to enhance the semantic context of stay points, a crucial step in regions like Egypt where linguistic diversity adds complexity to POI interpretation.
C.2.2 Transfer Learning and Model Training
Table 6: Model evolution during iterative transfer learning
JSD Values Activity Type Start Time End Time
NHTS vs Egypt GPS (baseline) 0.0255 0.0637 0.0433
Iteration_1 0.0055 0.0052 0.0027
Iteration_2 0.0038 0.0035 0.0021
Iteration_3 0.0039 0.0025 0.0021
Iteration_4 0.0027 0.0034 0.0029
Iteration_5 0.0043 0.0042 0.0045
Iteration_6 0.0029 0.0049 0.0046
To adapt the base model trained on the NHTS dataset, we employed transfer learning, leveraging the semi-supervised iterative refinement process detailed in our previous study. The adaptation process retained universal temporal patterns while progressively aligning with the distinct cultural and regional mobility characteristics of Egypt.
The progressive unfreezing approach was instrumental in balancing the preservation of global patterns with the learning of localized nuances. Initially, only the MLP and embedding layers were fine-tuned, allowing for rapid adaptation of activity classifications and temporal representations. In subsequent iterations, the transformer layers were gradually unfrozen to refine the model’s understanding of intermediate and complex activity dependencies. Its evolution through the semi-supervised transfer learning using Egypt GPS data is visualized and quantified in Figure 15 and Table 6.
Figure 15: Iterative adaptation from NHTS to Egypt via semi-supervised transfer learning with Egypt GPS data
The Egypt TUS data was aggregated to derive target activity duration distributions across the six demographic groups. Adapter training followed the same computational configuration as the Los Angeles case, utilizing a single NVIDIA RTX 4090 GPU. Training each demographic group required approximately 40 minutes, demonstrating the scalability of the approach.
The inference process was conducted on NVIDIA L40S GPUs, generating synthetic activity chains that closely aligned with the target distributions. As shown in Table 7, the JSD values for all demographic groups were low (ranging from 0.0054 for mid-aged males to 0.0302 for mid-aged females), indicating strong alignment between the generated and target distributions. This underscores the framework’s ability to replicate realistic activity patterns even in a data-scarce context.
Table 7: Egypt major activity duration (minutes/day) comparison with JSD
Demographic Group (JSD) Activity Duration (minutes)
Home Work School Buy goods Buy meals Recreational Exercise Visit friends
(1) (2) (3) (5) (7) (9) (10) (11)
Young Male (0.0067) 842.24/790.56 94.98/90.72 347.63/308.16 4.22/15.84 15.04/1.44 33.52/38.88 42.96/31.68 32.41/56.16
Young Female (0.0172) 786.10/761.76 76.75/70.56 316.58/313.92 3.13/53.28 14.41/4.32 131.45/148.32 1.43/8.64 36.24/53.28
Mid-Aged Male (0.0054) 684.13/688.32 342.37/348.48 12.70/10.08 8.87/14.40 15.00/15.84 136.65/126.72 4.07/4.32 119.86/112.32
Mid-Aged Female (0.0302) 881.72/832.32 199.05/146.88 63.73/0.00 8.07/24.48 19.69/5.76 146.69/126.72 0.33/4.32 36.20/53.28
Elderly Male (0.0141) 752.65/764.64 80.22/56.16 2.16/2.88 4.57/18.72 17.38/47.52 186.85/197.28 1.26/5.76 135.41/108.00
Elderly Female (0.0198) 848.55/853.92 107.94/79.20 0.00/17.28 3.10/18.72 10.85/4.32 164.29/142.56 0.00/11.52 46.59/106.56
Note: Values shown as TUS/Output. Each demographic group’s JSD value indicates overall distribution similarity (lower is better).
C.2.3 Observations on Generated Patterns
The generated activity chains demonstrated high consistency with target distributions, particularly for mandatory activities. Mid-aged males exhibited significantly longer work durations compared to mid-aged females, reflecting the disparity in employment rates (65% for males vs. 18% for females, as reported by the UN). Work durations for females in Egypt were notably lower than those in Los Angeles, illustrating the framework’s sensitivity to cultural and socio-economic differences. Religious Activities: Prominent across all demographic groups, with peaks reflecting the significance of prayer times in Egyptian society. Recreational Activities: Observed to be less frequent, consistent with the conservative cultural context of Egypt compared to Western regions like Los Angeles.
The Egypt case study demonstrates the framework’s adaptability to data-scarce regions, effectively generating realistic activity chains aligned with local sociodemographic patterns and cultural contexts. By leveraging transfer learning, adapter training, and LLM-based semantic enrichment, the framework overcomes the challenges of sparse data and fragmented sources, showcasing its potential for mobility modeling in diverse global settings.